### v1

# Improving Multi-candidate Speculative Decoding

Abstract
Speculative Decoding (SD) is a technique to accelerate the inference of Large Language Models (LLMs) by using a lower complexity draft model to propose candidate tokens verified by a larger target model. To further improve efficiency, Multi-Candidate Speculative Decoding (MCSD) improves upon this by sampling multiple candidate tokens from the draft model at each step and verifying them in parallel, thus increasing the chances of accepting a token and reducing generation time. Existing MCSD methods rely on the draft model to initialize the multi-candidate sequences and use static length and tree attention structure for draft generation. However, such an approach suffers from the draft and target model’s output distribution differences, especially in dynamic generation context. In this work, we introduce an improved version of MCSD that includes a target model initialized multi-candidate process, dynamic sliced topology-aware causal mask for dynamic length adjustment, and decision models to optimize early stopping. Our framework improves the acceptance rate $\alpha$, defined as the ratio of the longest draft sequence length accepted by the target model over the maximum draft sequence length, by a maximum of 164% and gains a maximum of 75% generation speed up over the MCSD baseline. We also conduct an ablation study to evaluate the impact of the decision model.
1 Introduction
In recent years, Large Language Models (LLMs) such as GPT-4achiam2023gpt have significantly advanced various language processing tasks. However, these models are computationally intensive especially during the inference phase, where generating $k$ tokens requires $k$ serial runs of the model. This inefficiency limits the practical deployment of these powerful models in real-time applications.
Among serious LLM inference optimization methods, Speculative Decoding (SD)leviathan2023fast has been shown to increase inference speed with a marginal generation quality loss. SD frameworks first generate candidate tokens using a model (the draft model) with lower complexity compared to the original model (the target model). Then the target model verifies the generated tokens. The performance of SD is mainly determined by the token acceptance rate $\alpha$, which measures the proportion of candidate tokens generated by the draft model that the target model accepts. The benefits of using a faster draft model diminish if the target model frequently rejects these tokens, which means that the target model must re-generate the tokens.
To further enhance the acceptance rate, Multi-Candidate Speculative Decoding (MCSD)10.1145/3620666.3651335 ; yang2024multi was introduced. MCSD samples multiple candidate tokens at each generation step and verifies them in parallel using the target model. This approach increases the likelihood that at least one of the candidate tokens will be accepted, thereby improving the overall acceptance rate and efficiency. MCSD also incorporates a tree attention mechanism to manage computational and communication overhead by organizing multiple candidate sequences into a single sequence and applying a carefully designed attention mask. However, MCSD still faces several challenges. 1. Increasing Computational Complexity: Verifying multiple candidates simultaneously increases the computational load, requiring more memory and processing power. 2. Efficient Topology-Aware Causal Mask Generation: Generating and maintaining a topology-aware causal mask for multi-candidate token trees is time-consuming and reduces the adaptivity of the model. 3. Fixed Draft Generation Length ($\gamma$): Using a fixed length for draft-generated token segments may not be optimal in all contexts.
In this paper, we present a method that introduces dynamic sliced topology-aware causal mask to facilitate the speculative decoding process, allowing dynamic adjustment of the draft generation length without reconstructing the topology-aware causal mask. We enhance the acceptance rate by initializing the multi-candidate token tree with the target model, thus improving efficiency. Additionally, we incorporate a decision model to optimize the early stopping mechanism during the draft model generation stage. The model dynamically halts draft token generation early by predicting the likelihood of the target model accepting the tokens, thus reducing unnecessary computation.
Our experimental results show that our framework significantly improves the acceptance rate ($\alpha$), defined as the ratio of the longest draft sequence length accepted by the target model to the maximum draft sequence length. On average, it achieves a 40% increase in acceptance rate and a 23% speedup in generation time compared to the MCSD baseline yang2024multi on the larger draft model (TinyLlama-1.1Bzhang2024tinyllama ). These improvements are even more significant when using the smaller draft model (LLaMa-68M10.1145/3620666.3651335 ), with increases of 164% in acceptance rate and 75% in generation speed. All improvements are measured using prompts from MT-Benchzheng2024judging . We also conduct an ablation study to evaluate the impact of the decision model. Our results show that the target model initialized multi-candidate token tree contributes the most to speed up gains. While the decision models we explored only speed up MCSD when the width of the token tree is small enough. Future dynamic multi-candidate speculative decoding work may benefit from the dynamic sliced topology-aware causal mask to reduce the overhead caused by casual mask for each draft generation iteration.
2 Background
Figure 1: $\beta$ denotes the threshold for early stop, in our experiment the $\beta$ = 0.4. The inputs for the decision model are hidden states or output distribution and entropy related to the token. The dynamic multi-candidate speculative decoding process with an early stop decision model and fork-shaped draft model initialized a token tree, where the token tree configuration is width=3 and depth=3. It stopped at the second draft generation turn, where the maximum number of draft generation turns is three.
2.1 Speculative Decoding
Speculative decoding is the collaboration of two models: a smaller draft model (often a more efficient approximation model) and a larger target model. First, the draft model generates multiple candidate tokens in parallel, using its probability distribution to predict the following possible tokens based on input. Then, these tokens are passed to the target model, which verifies them by computing its probability distribution over the same input. If the target model accepts the candidate tokens (probabilities alignment), they are finalized in the output sequence. Otherwise, the target model replaces them by generating new tokens based on its own distribution. The speculative decoding process ensures that the output distribution remains consistent with what the target model alone would produce, thus maintaining the quality of the generated contentleviathan2023fast . Importantly, this technique does not require changes to the model’s architecture or retraining, making it an accessible and efficient solution for accelerating inference.
2.2 Multi-candidate generation
Due to distributional differences between the draft and target models, the candidate path with the highest probability in draft model may not always result in the most accepted tokens by the target model. Therefore, verifying multiple candidate paths in parallel increases the overall acceptance rate $\alpha$ of draft tokens. This paper takes the tree attention methodology in Specinfer 10.1145/3620666.3651335 as a starting point to process multiple candidate token paths concurrently. In Figure 2, we illustrate the topology-aware causal mask for calculating the tree attention of a multi-candidate sequence with three candidate token paths, each with two draft tokens.
3 Method
Figure 2: Both draft-initialized and target-initialized multi-candidate generation processes use width=3 and depth=2 as token tree configurations.
In the original speculative decoding framework leviathan2023fast , the expected improvement factor (IF) is defined as:
$IF=\frac{1-\alpha^{\gamma+1}}{(1-\alpha)(c\gamma+1)}$
where $\alpha$ represents the expectation of acceptance rate, $\gamma$ represent the draft generation length, and $c$ represent the ratio between the time for a single run of draft model and time for a single run of target model. Based on the formula (1), a larger $\alpha$ will speed up token generation. If an oracle could determine $\gamma$ dynamically, the improvement factor can be up to around 60% larger than the improvement factor with a fixed $\gamma$.
In this work, we explore three methods to improve $\alpha$ and determine $\gamma$ dynamically to accelerate the MCSD process:
  1. 1.
    Target Initialized Multi-Candidate Token Tree: we introduce a new methodology to construct multi-candidate sequences that improve $\alpha$ over existing approaches;
  2. 2.
    Dynamic Sliced Topology-Aware Causal Mask: we introduce a method to efficiently creating topology-aware casual masks for dynamic multi-candidate generation;
  3. 3.
    Early Stop Decision Model: we introduce a low-complexity MLP model to determine $\gamma$ dynamically during each draft generation loop.
We remark that the above improvements can be integrated into a unified framework. The Target Initialized Multi-Candidate Token Tree method can be deployed in isolation as it improves $\alpha$ for both static and dynamic MCSD, while $2$ and $3$ are dependent on each other and need the dynamic MCSD.
3.1 Target Model initialized multi-candidate token tree
Existing multi-candidate speculative decoding methods rely on a draft model to generate the entire multi-candidate token tree, and after verifying the draft generated tokens, only sample one token from the target model or normalized target and draft model’s output distribution. Due to the difference between the output distribution of the draft and target model, there is no specific criterion to determine which token sampled from the target model will yield the longest accepted token sequence for future draft model generation. Therefore, we hypothesize that sampling multiple tokens instead of one token from the target model’s distribution to initialize a multi-candidate sequence for future draft model generation can increase the acceptance rate.
Figure 2 illustrates how we modified the existing draft-initialized multi-candidate generation to build our target model-initialized multi-candidate generation. The topology-aware mask for the target-initialized token tree needs a larger mask than the draft-initialized token tree to handle the multiple initial tokens for different possible sequences sampled from the target model. The topology-aware causal mask size for the target initialized token tree is the square of the sum of the token tree width and the total number of draft-generated tokens. In Figure 2, the size of the mask for the target initialized token tree is $(3+6)^{2}$ while the size of the mask for the draft initialized token tree is $6^{2}$. We note that if draft-initialized and target-initialized multi-candidate generation has the same acceptance rate, then the target-initialized method will perform one more target forward pass compared to the draft-initialized method as the origin input passes into the target model instead of the draft model. However, in general, text generation tasks usually involve hundreds of target forward passes, and this overhead is marginal.
In our setup, we use LLaMA-68M10.1145/3620666.3651335 as a draft model and LLaMA 2-7B Chattouvron2023llama as the target model. Figure 3 shows that using the target model to initialize a multi-candidate token tree increases the percentage for paths other than 0, yielding the longest acceptance rate, and increases the average accepted length for all paths. Thus, sampling multiple tokens from the target model’s distribution to initialize a multi-candidate sequence increases the acceptance rate.
3.2 Dynamic Sliced Topology-Aware Causal Mask
Existing multi-candidate speculative decoding methods such as EAGLE li2024eagle employ expansion-based token trees with various depths and widths for different branches to increase the target model’s average acceptance rate on draft-generated tokens. Generating a topology-aware casual mask for the multi-candidate token tree is time-consuming; most existing multi-candidate speculative decoding methods only build the topology-aware casual mask once during initialization.
We introduce a dynamic sliced topology-aware casual mask to allow the decision model to dynamically decide the length of multi-candidate draft token generation and avoid generating a new topology-aware casual mask during each iteration. The main idea is to construct a large topology-aware casual mask during initialization. Figure 1 illustrates a topology-aware casual mask that allows a maximum of three draft token generations for each candidate sequence. There are three possible sliced causal masks for dynamic multi-candidate speculative decoding: mask when early stop at first iteration (upper left 3 x 3 causal mask), early stop at second iteration (shown in Figure 1), no early stop (entire 9 x 9 casual mask).
In the proposed framework, we keep the token tree in fork shape so that each candidate token sequence does not expand to new sub-sequences. The main benefit of using a fork shape instead of an expansion-based token tree is a reduction in memory usage, since the size of a topology-aware causal mask for an expansion-based token tree grows exponentially. In contrast, the fork-shaped token tree will grow linearly.
Figure 3: Left graph shows the frequency of which path yields the longest accepted length. The right graph shows the average accepted length for each path; the maximum length is five. In this graph, the target and draft model initialized multi-candidate token tree has static six candidate paths and a depth of five for each path. Both methods run on the same set of inputs.
3.3 Decision Model
We design two types of decision models to dynamically determine if the early stop is necessary for the draft generation process. The first type of decision model is a three-layer MLP, taking the hidden states from draft model as input and the $max(1,\frac{p(x|I)}{q(x|I)})$ ($p(x|I)$ is the probability of target model predict token $x$ with given input $I$ and $q(x|I)$ is the probability of draft model predict token $x$ with given input $I$) as the training label. The decision process can be represented as:
$P_{T1}=MLP(y_{i}^{H})$
Where $y_{i}^{H}$ denotes the hidden states of the draft model for token $i$.
The second type of decision model is inspired by Tandem transformer s2024tandemtransformersinferenceefficient , is a two layer-MLP with draft model’s output distribution’s entropy that takes probabilities as input and the result of verification (zero as rejected, one as accepted) as a label. The decision process can be represented as:
$P_{T2}=MLP(ConCat(Topk(y_{i}^{D}),Ent(y_{i}^{D}))$
Where $y_{i}^{D}$ represents the output distribution of the draft model for token $i$ and is used to compute both the entropy and top-k probabilities for input to the two layer-MLP.
During the draft generation process, the decision model will batch inference the input for multiple sequences and calculate the probabilities of the target model accepting each sequence. If the probabilities for all sequences are lower than the threshold (in our experiment, the threshold is 0.4 for both decision models), then draft generation will stop early; otherwise, it will continue until the maximum draft generation length is reached.
4 Experimental Results
In this section, we analyze the performance of the methods we introduced in the previous sections.
4.1 Experimental Setup
Our experiments are conducted on a single server with an Nvidia RTX 4090 GPU. All target models are configured with Bfloat16 precision to match the Llama model’s original training weight, while draft models utilize double precision. This setup creates a stable testing environment and ensures that the difference in the inference speed of the models is based on the inference optimization method rather than the model accuracy and size.
For the draft and target models, we extensively utilize the LLaMA 2-7B Chat modeltouvron2023llama and other related models that share the same tokenizer. As for our target models, we select LLaMA 2-7B Chat and its fine-tuned version, Vicuna-7Bvicuna2023 . At the same time, we use two different models, the Llama-68M from SpecInfer 10.1145/3620666.3651335 and the TinyLlama-1.1Bzhang2024tinyllama , as draft models to test the effect of the acceleration methods on different types of model pairs. The maximum generation length is 200 tokens. We also employ temperature settings of 0 (greedy sampling) and 1 (probabilistic sampling).
Furthermore, we standardize the configuration across different SD methods to evaluate the acceptance rate and other performance metrics. Specifically, we set the SD to a fixed $\gamma=5$, while for the MCSD method, we use the optimal k-configuration (4,2,2,1,1) from the original paper yang2024multi . Our method employs a dynamic configuration with $D=5$ and $W=16$, where $W$ (Width) indicates the number of candidate sequences generated in parallel at each step of the token tree, and $D$ (Depth) represents the level of token prediction, referring to $\gamma$ in traditional SD. The fixed $D=5$ ensures the draft generation length is consistent across all SD methods. At the same time, $W=16$ aligns the number of candidate sequences with those in MCSD, enabling a fair comparison between the methods.
For our testing dataset, we select TriviaQAjoshi-etal-2017-triviaqa , Alpacaalpaca and MT-Benchzheng2024judging . In the following testing, for TriviaQA and Alpaca, we randomly select 250 input prompts. And for MT-bench, we test the complete 80 sets of prompts.
Dataset Draft models Temp Acceptance Rate $\alpha$
LLaMA 2-7B Chat Vicuna-7B
Alpaca LLaMA-68M 0 0.80 0.83
TinyLlama-1.1B 0 0.92 0.94
LLaMA-68M 1 0.75 0.79
TinyLlama-1.1B 1 0.93 0.95
TriviaQA LLaMA-68M 0 0.86 0.91
TinyLlama-1.1B 0 0.96 0.94
LLaMA-68M 1 0.75 0.80
TinyLlama-1.1B 1 0.93 0.94
MT-Bench LLaMA-68M 0 0.76 0.77
TinyLlama-1.1B 0 0.93 0.93
LLaMA-68M 1 0.74 0.75
TinyLlama-1.1B 1 0.95 0.95
Table 1: Acceptance rates of our methods (temp = 0 and 1, given generation depth $D=5$ and width $W=16$)
4.2 Overall Results
Figure 4: Speedup ratios compared to vanilla inference for different SD methods based on MT-Bench dataset using greedy sampling. The bars represent speedup ratios for different model combinations (draft model + target model)
Figure 5: Left graph shows the relationship between width and acceptance rate ($\alpha$) and right graph shows the the relationship between width and generation speed in tokens/s
Dataset Methods Configuration LLaMA 2-7B Chat Vicuna-7B
Alpaca Baseline SD $\gamma=5$ 0.45 0.43
MCSD $4\times 2\times 2\times 1\times 1$ 0.64 0.63
Our method $D=5,W=16$ 0.92 0.94
TriviaQA Baseline SD $\gamma=5$ 0.56 0.42
MCSD $4\times 2\times 2\times 1\times 1$ 0.72 0.59
Our method $D=5,W=16$ 0.96 0.94
MT-Bench Baseline SD $\gamma=5$ 0.43 0.49
MCSD $4\times 2\times 2\times 1\times 1$ 0.63 0.68
Our method $D=5,W=16$ 0.93 0.93
Table 2: Comparison of acceptance rate ($\alpha$) for different methods using TinyLlama-1.1B as draft model with greedy sampling (temp = 0)
The experiment result shown in Figure 4 demonstrates substantial improvements in generation speed. All of the configurations are shown and explained in Section 4.1. Conducted on the MT-Bench dataset, our method achieves a maximum of 2.96 times speedup over the baseline, mainly due to the target model initialized multi-candidate token tree.
Also, we measure the effect of width on acceptance rate and speed in Figure 5. As the width increases, we consistently observe an improvement in acceptance rates and speed across different model pairs. The acceptance rate curve converges when width=12, and the speed curve converges in width=10. These findings demonstrate our method’s effectiveness in improving acceptance rates and speed with relatively small tree widths. In addition, we notice there is a noticeable drop in speed when width=21. This speed decline is due to the increased computational overhead associated with processing a more significant number of candidate tokens in parallel, which begins to outweigh the benefits of speculative decoding.
4.3 Acceptance Rate
We run acceptance rate $\alpha$ test across all three datasets with two different temperatures, shown in Table 1. Table 2 shows our method consistently demonstrating higher acceptance rate $\alpha$ across various datasets compared to baseline SD and MCSD. This indicates that our method’s dynamic depth adjustment and target model initialization significantly enhance its ability to generate sequences more aligned with the target model’s expectations. This leads to a higher overall acceptance rate, even under the same draft generation constraints.
4.4 Ablation Study
Figure 6: Both graphs show the relationship between width and generation speed. Draft model on the left graph: Llama-68M. Draft model on the right graph: TinyLlama-1.1B
To show the effect of the decision model in our framework, we compare the inference speeds under two scenarios, using or not using the decision model with different parameter size draft models. We still use MT-bench as input prompts, and all other settings remain identical. As a result, in Figure 6, both large and small draft models return higher speeds without using the decision model most of the time. The decision model only improves the speed slightly when the width is low. However, in either case, the speed improvement of the decision model is minimal. In addition, it proves that our method’s speedup is mainly due to the target model’s initialized multi-candidate token tree and topology-aware causal mask rather than the decision model.
5 Conclusions
In this work, we proposed a target model initialized multi-candidate token tree that successfully increases the acceptance rate for multi-candidate speculative decoding and a dynamic mask slicing technique to avoid topology-aware causal mask generation overhead for dynamic multi-candidate speculative decoding. While we have yet to discover the decision model that could make dynamic multi-candidate speculative decoding faster than static one under all scenarios, our work can help future research to find a better decision model and process for dynamic speculative decoding.